killerjay556:

Given three side lengths, select the set for which a triangle exists. Group of answer choices 6 cm, 14 cm, 8 cm 8.5 cm, 17 cm, 10.6 cm 9 cm, 11 cm, 21 cm 14 cm, 4.7 cm, 4.7 cm

itsmehjay:

Are we talking about a specific triangle like a right triangle? If so then we need to apply the Pythagorean Theorem -> a^2 + b^2 = c^2

NerdyLizard:

What I said and then deleted since it did not say right angle

Florisalreadytaken:

no well if the order of the lengths of the legs is leg, leg, hypotenuse then yeah only of the options is right the sum of the adjacent and opposite leg is greater than the length of the hypotenuse -- basically, just try every single one of them, and whichever turn to be right, that's your answer.

NerdyLizard:

Lol forgot about that rule thanks @florisalreadytaken

itsmehjay:

Also remember, the numbers have to be in order so least to greatest

NerdyLizard:

I am unsure if that is true.

itsmehjay:

Yeah they have to be or else it will be incorrect

Florisalreadytaken:

@itsmehjay wrote:
Also remember, the numbers have to be in order so least to greatest
what?

NerdyLizard:

None of them are labeled. They can be switched.

Florisalreadytaken:

as i mentioned, if the order is ` leg, leg, hypotenuse ` then only 1 of the options is wrong -- just try them e.g lets try a) 6 cm, 14 cm, 8 cm \( 6+16>8 \) that is true so that works -- -- -- -- -- -- -- -- 8.5 cm, 17 cm, 10.6 cm \(8.5+17>10.6\) that is true as well -- -- -- -- -- -- -- -- 9 cm, 11 cm, 21 cm \(9+11>19\) i dont think that works 14 cm, 4.7 cm, 4.7 cm \(14+4.7>4.7\) obviously, that works as well but if the order is stated to be different please mention

surjithayer:

if a,b,c are the sides of a triangle,then 1. sum of any two sides >third side. 2. difference of any two sides <third triangle.

surjithayer:

only B satisfies these properties. 17-8.5=8.5<10.6

surjithayer:

10.6+8.5=19.1>10.6

Florisalreadytaken:

@surjithayer wrote:
2. difference of any two sides <third triangle.
difference of any two sides < third side* and yes, that was for killerjay556 to mention -- logically if the sum of the two is greater, then the difference would be smaller than the 3rd side.

surjithayer:

14-6=8 not less than third side it is a triangle of zero height . or we can say it is not a triangle.

surjithayer:

mostly sum of two sides > third but difference of two sides must be < third side for it to be a triangle.